OpenStudy (lilcathyx):

ok i really don't understand this :O Help ? 3. In step 1, you found the volume (in cubic feet) of the main tank. If the maximum density of killer whales per cubic foot is 0.000011142, what is the maximum number of killer whales allowed in the main show tank at any given time?

OpenStudy (lilcathyx):

Volume of tank #1 is 359007 cubic feet.

OpenStudy (lilcathyx):

@jim_thompson5910 if it doesnt bother you, could u help me?

jimthompson5910 (jim_thompson5910):

(volume of tank)*(density of killer whales) = number of killer whales allowed (359007 cubic feet)*(0.000011142 whales/1 cubic foot) = 359007*0.000011142 = 4.000055994 so you're allowed to have 4 killer whales

jimthompson5910 (jim_thompson5910):

I multiplied because the goal is to get the units "cubic feet" to cancel leaving the number of whales as the only unit

jimthompson5910 (jim_thompson5910):

4 killer whales sounds like a reasonable amount (when thinking of any theme park) it might be high and 3 might be better. I'm not sure as I'm not a biologist

OpenStudy (lilcathyx):

so u divided by 1 right ?

OpenStudy (anonymous):

Yes to cancel the units

jimthompson5910 (jim_thompson5910):

yeah 1 cubic foot

jimthompson5910 (jim_thompson5910):

density = 0.000011142 whales per cubic foot = (0.000011142 whales)/(1 cubic foot)

OpenStudy (anonymous):

When you cancel the units, the only unit left is kill whales

OpenStudy (anonymous):

Which is what you are looking for. You want to know the number of killer whales

OpenStudy (lilcathyx):

i see it what u mean and how Jim did it. so this is already in cubic feet?

jimthompson5910 (jim_thompson5910):

what is already in cubic feet?

OpenStudy (lilcathyx):

4.000055994 = 4 4 killer whales right there's no cubic feet here.

OpenStudy (anonymous):

Yup 4 killer whales. Look at Jims first reply. It explains it

OpenStudy (lilcathyx):

Well thank you both ! :D :)^.^ i understand now how to do this for tomorrows test ! :D

jimthompson5910 (jim_thompson5910):

yeah the units for the answer is the number of whales you're welcome